Three urns of same appearance have balls
Urn1: 2 black and 1 white
Urn 2: 1 black and 2 white
Urn 3: 2 black and 2 white
If a urn is selected and one ball drawn is white what is the probability that second ball drawn is white if the first one is not returned
 

Grade:12th pass

1 Answers

Saurabh Koranglekar
4 years ago
Dear student

Selection of white from urn 2 = 0.33 * 0.5

Selection of white from urn 3 = 0.33 *0.3

Required probability = 0.33* 0.8 = 0.264

Regards